Write a quadratic equation in the variable x having the given numbers as solutions. Type the equation in standard form, ax^2+bx+c=0. Solution: 4,only solution.
Thank you

Answer by stanbon(75887)


Write a quadratic equation in the variable x having the given numbers as solutions. Type the equation in standard form, ax^2+bx+c=0. Solution: 4,only solution.
--------------
If x = 4 is the only solution, it must have multiplicity 2.
---
Ans: (x-4)^2 = 0
x^2-8x+16 = 0
=====================
Cheers,
Stan H.
===============